OpenStudy (anonymous):

can someone help me. I don't want the answer I want someone to walk me through it. Given that f(x) = 6x + 2 and g(x) =(2x+4)/5 for g(f(1)).

OpenStudy (anonymous):

4

OpenStudy (anonymous):

haha ok thanks. I didn't want the aswer but I guess that works:)

OpenStudy (anonymous):

\[g(f(x))= (2(6x + 2)+4))/5 g(f(1))=(2(6.1 + 2)+4))/5 =20/5 =4\]

OpenStudy (anonymous):

could you help with this one? Given f(x) = (3x-1)/2 solve for f-1(4). @boyb39 @rav9027

OpenStudy (anonymous):

ask question clearly